A substance is dissolved in water,forming a 0.5 molar solution.If 4.0L of the solution contains 240g of the substance,what is the molecular mass of the substance ?

Grade:8

2 Answers

Arun
5 years ago
M = m/MM x V, where m = mass, MM = molar mass, V = volume in liters 
MM = m/MxV ---> 240-g/0.50-mol/L x 4-L = 120 g/mol 
Ans = 120 is molecular mass of unknown
Jc Chaudhary
5 years ago
Simple,
Molarity = no. Of moles /volume 
Molarity = mass given /molecular mass/volume
0.5=240/mm/4
1/mm=1/120
MM = 120g/mol